OpenStudy (noahjohnson967):

Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. -185 -77 -23 85

OpenStudy (anonymous):

Are the 3 after the x, 2 after y and 4 after z supposed to represent powers?

OpenStudy (noahjohnson967):

yes sorry

OpenStudy (noahjohnson967):

its the last problem on the last assignment i have to do today and im just stuck

OpenStudy (anonymous):

just plug 3 in for x, 5 in for y and -3 in for z then solve. x^3=3^3=27

OpenStudy (noahjohnson967):

i keep getting answers that arent even an option

OpenStudy (mathstudent55):

\(x^3 - 2y^2 - 3x^3 + z^4\) Is this the correct expression?

OpenStudy (noahjohnson967):

yes

OpenStudy (mathstudent55):

The first step you can do is combine like terms. The first term is x^3 and the third term is -3x^3. Those terms are like terms, so they can be combined. What is x^3 - 3x^3 = ?

OpenStudy (mathstudent55):

Are you still there?